# Black-box Membership Inference Attacks against Fine-tuned Diffusion Models

### v1

@@ Abstract
With the rapid advancement of diffusion-based image-generative models, the quality of generated images has become increasingly photorealistic. Moreover, with the release of high-quality pre-trained image-generative models, a growing number of users are downloading these pre-trained models to fine-tune them with downstream datasets for various image-generation tasks. However, employing such powerful pre-trained models in downstream tasks presents significant privacy leakage risks. In this paper, we propose the first reconstruction-based membership inference attack framework [^1] Code accessible at https://anonymous.4open.science/r/Reconstruction-Attack, tailored for recent diffusion models, and in the more stringent black-box access setting. Considering four distinct attack scenarios and three types of attacks, this framework is capable of targeting any popular conditional generator model, achieving high precision, evidenced by an impressive AUC of $0.95$.
1 Introduction
The recent developments in image-generative models have been remarkably swift, and many applications based on these models have appeared. Diffusion models [41, 18, 35, 40, 46, 53, 56, 44, 19, 54] have come to the forefront of the image generation. These models generate target images by progressive denoising a noisy sample from isotropic Gaussian distribution. In an effort to expedite the training of diffusion models and reduce training expenses, Stable Diffusion [44] was introduced. Leveraging the extensive and high-fidelity LAION-2B [48] dataset for training, the Stable Diffusion pre-trained checkpoint, available on HuggingFace, can be fine-tuned efficiently with just a few steps for effective deployment in downstream tasks. This model’s efficiency has spurred an increasing number of usages of Stable Diffusion.
@@ 1 Introduction
In a white-box attack scenario, attackers have access to all parameters of a model. Similar to membership inference attack targeting classifiers, attacks against diffusion models also utilize internal model information such as loss [6, 20, 34] or gradients [37] as attack features. Although white-box attacks can achieve high success rates, their limitation lies in the requirement for complete access to the target model’s information, which is often impractical in real-world scenarios. Compared with white-box attack, gray-box approaches do not require full access to the model’s parameters; instead, they only necessitate the intermediate outputs from the diffusion model during the denoising process to serve as features for inference [12, 14, 20, 25]. However, these attacks inevitably rely on the intermediate images generated during the model’s operation. In real-world scenarios, if a malicious model is trained using private or unsafe images, typically only the final output image is provided, with efforts made to conceal as many model details as possible. Therefore, the more practical scenarios would be black-box.
@@
In the black-box setting, there are attacks for GANs [7, 17] and VAEs [17]. These are based on unconditional generative models and involve a highly stochastic generation process that requires extensive sampling for inference, which becomes inefficient when directly applied to diffusion models. The other black-box attacks [61, 63], although more tailored for diffusion models, focus on simulations and lack the necessary conditions to be used in realistic scenarios. We will discuss them in Section 2.4.
@@
In this paper, we present a novel black-box attack framework fit for state-of-the-art image-generative models. The framework was built on a careful analysis of the objective function of the diffusion model as its theoretical foundation. It also incorporates four potential attack scenarios tailored for different settings of diffusion models. We demonstrate the efficacy of our attack using a Stable Diffusion model and further validate it on models fine-tuned with CelebA-Dialog [23], WIT [57], and MS COCO datasets [30].
@@
  • Consistent with the definition in Suya et al. [58], four attack scenarios are considered in which an attacker can perform an attack based on the query access as well as the quality of the initial auxiliary data, and three different inference models are used to determine the success rate of the attack, respectively.
  • The efficacy of the attack is evaluated on the CelebA, WIT, and MS COCO datasets using Stable Diffusion as the representative target model. The attack’s impact is analyzed by considering various factors: image encoder selection, distance metrics, fine-tuning steps, inference step count, member set size, shadow model selection, and the elimination of fine-tuning in the captioning model.
@@
Section 2 reviews key works on denoising generative models and membership inference attacks, including their application against diffusion models. Section 3 introduces our reconstruction-based black-box attack on diffusion models, tailored to four levels of attacker knowledge. Section 4 describes our experimental setup, and Section 5 compares our attack’s effectiveness with existing methods and examines various influencing factors.  Section 6 shows that our attacks remain effective under common defenses.  Section 7 discusses some other research related to our work.  Section 8 concludes the paper, summarizing our main findings and contributions.
2 Background
2.1 Machine Learning
@@
2.1.1 Classification Models
In the context of classification model training, the objective is to map an input $x$ to a category $y$. The functional representation of the model can be expressed as $y=\mathcal{M}(x)$, where $x$ denotes the input image, and $y$ signifies the corresponding label. The loss in a classification model, which quantifies the discrepancy between the predicted and true labels, can be articulated as follows:
@@ 2.1.1 Classification Models
2.1.2 Generative Models
Generative models are designed to generate images $\hat{x}=\mathcal{M}(z)$, where $z$ is the randomness not provided by users by inherent to the server hosting the model.
Generative models include VAEs [24], GANs [15], and diffusion models [18]. Recently, diffusion models have gained significant traction. Building on the classical DDPM (Denoising Diffusion Probabilistic Models), a plethora of models, such as Imagen [46], DALL·E 3 [3], GLIDE [35], Stable Diffusion [44], have emerged and can generate high-quality images based on prompt information. In this paper, we focus on diffusion models.
Fundation of Diffusion Models.
@@ 2.1.2 Generative Models
In this equation, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, where $\alpha_{i}$ is a predefined parameter that decreases incrementally within the interval $[0,1]$. The term $\epsilon_{t}$ is a random Gaussian noise derived using the reparameterization trick from multiple previous forward steps (more details in Appendix A).
@@
Prompt Guided Diffusion Models.
While various models [3, 40, 46, 35, 44] mentioned above are capable of generating high-quality images conditional on prompt information $p$, describe as $\hat{x}=M(z,p)$, further details can be found in Appendix B. Our experiments primarily utilize the current publicly available state-of-the-art model, Stable Diffusion [44]. Distinct from other diffusion generative models [35, 40, 46], Stable Diffusion uniquely conducts both the forward and reverse processes within the latent space. This approach offers advantages: the noise addition and removal processes operate over a smaller dimensionality, allowing for faster model training at lower computing costs. Additionally, within the latent space, the model can accommodate diverse prompt information to guide image generation. Importantly, Stable Diffusion is open-sourced and provides multiple high-quality pre-trained checkpoints online. This aligns well with the focus of our study on potential privacy concerns when fine-tuning pre-trained models for downstream tasks.
2.2 Membership Inference Attacks
@@
Carlini et al. [5] suggest the use of confidence scores as attack features and the creation of two distributions, $\mathbb{D}_{\text{in}}$ and $\mathbb{D}_{\text{out}}$, based on the confidence scores of samples from the member and non-member sets, respectively. The distributions are then utilized to calculate the probability density function of a query data point $x$ in the member set and non-member set.
@@
MIAs against Diffusion Models
@@
In the context of MIA against diffusion models, due to the structural differences between diffusion models and classification models, as well as the dissimilarities in their inputs and outputs, MIAs designed for classification models cannot be directly applied to diffusion models. The focus of research lies in how to construct features for MIA. We classify existing attacks against diffusion models as white-box, gray-box, and black-box, and introduce them separately. In white-box attacks, methods in this setting exploit the loss (derived from each timestep using Equation 2) and gradients (via backpropagation through the model). Gray-box attacks typically necessitate access to a model’s intermediate outputs but do not require any internal model information. For gray-box attacks targeting diffusion models, the model’s denoising trajectory, particularly the noisy images, is utilized as attack data. In contrast, black-box attacks operate without knowledge of the model’s internal mechanics or process outputs, relying solely on the final generated images for analysis. In Table I, we compare all existing attacks. Each type of attack’s details are deferred to Section 2.4.1 and Section 7.
2.3 Problem Formulation
2.3.1 Threat Model
@@
Given the target sample $x$ and black-box access to the target image-generative model $\mathcal{M}$, the goal of the attacker is to determine whether $x$ was used to train $\mathcal{M}$. More specifically, we focus on the fine-tuning process, namely, we care about the privacy of the fine-tuning dataset of $\mathcal{M}$, and do not care about the pre-training dataset. We focus on fine-tuning because (1) the attacks will be similar for direct training, while the computational cost for experiments on fine-tuning MIA will be much smaller, and (2) the pertaining and fine-tuning paradigm is more popular with modern large models, and we will provide motivations for focusing on the fine-tuning process later.
@@
We categorize the threat model into four scenarios with two dimensions, namely:
@@
  • Target Sample Component. One distinct property of the current image generator models is that there exists the flexibility to input text prompt or initial images to guide model generation. Two configurations for the attacker can be considered: First, the query data $x$ aligns with the training data as a text-image pair ($x=\langle T_{q},I_{q}\rangle$, where $T_{q}$ denotes the text component and $I_{q}$ denotes the corresponding image component). Second, the attacker only obtains a suspect image potentially revealing private information without a corresponding caption ($x=I_{q}$). As our focus here is MIAs on text-image generative models, the scenario where $x$ solely consists of text is not deemed practical and hence, is not discussed.
@@
  • Auxiliary Dataset. Similar to all other MIAs, we assume an auxiliary dataset $\mathbf{D}^{\prime}$ is available. It is used to train the shadow models $\mathcal{M}^{s}$ to mimic the behavior of the target model. We consider two scenarios for $\mathcal{D^{\prime}}$: $\mathbf{D}^{p}$ and $\mathbf{D}^{s}$, indicating whether the attacker has access to real samples used to fine-tune $\mathcal{M}$ or just samples from the same distribution.
@@
2.3.2 Motivation
For a malicious publisher, training a generative model from scratch to produce malevolent images based on prompt information would entail significant computational costs. Therefore, the most efficient strategy for them would be to fine-tune a model that has already been pre-trained on a large-scale dataset. Given the full open-source nature of the Stable Diffusion [44], and the extensive availability of pre-trained models capable of generating photorealistic images from entities like CompVis [^2] https://huggingface.co/CompVis/stable-diffusion and Stability AI [^3] https://github.com/Stability-AI/generative-models, there has been an increasing trend of leveraging these pre-trained models for fine-tuning to specific downstream tasks. Furthermore, an increasing number of companies, such as Amazon [^4] https://aws.amazon.com/sagemaker/jumpstart/, OctoML [^5] https://octoml.ai/blog/the-beginners-guide-to-fine-tuning-stable-diffusion/, and CoreWeave [^6] https://docs.coreweave.com/cloud-tools/argo, are offering services in this domain. The data privacy issue during the training of these downstream tasks has not been explicitly studied. Our work seeks to uncover privacy leakage issues in this process and raise awareness of them.
@@ 2.3.2 Motivation
2.4 Existing Solutions
2.4.1 Black-box MIA against Traditional Image-generatve Models
@@
GAN-Leaks [7] employs a similar intuition and, instead of density in the Monte-Carlo attack [17]), roughly speaking, use the shortest distance of the generated samples from the target sample as the criteria. It also proposes another attack assuming an extra ability to optimize the noise input $z$ to the generator (which is not strictly the black-box setting; we will describe it and compare with it the evaluation) so it can reduce the number of generated samples. More formal details about these attacks are deferred to Appendix C.
2.4.2 Black-box MIA against Recent Diffusion Models

### v4

@@ Abstract
With the rapid advancement of diffusion-based image-generative models, the quality of generated images has become increasingly photorealistic. Moreover, with the release of high-quality pre-trained image-generative models, a growing number of users are downloading these pre-trained models to fine-tune them with downstream datasets for various image-generation tasks. However, employing such powerful pre-trained models in downstream tasks presents significant privacy leakage risks. In this paper, we propose the first scores-based membership inference attack framework [^1] Code accessible at https://github.com/py85252876/Reconstruction-based-Attack tailored for recent diffusion models, and in the more stringent black-box access setting. Considering four distinct attack scenarios and three types of attacks, this framework is capable of targeting any popular conditional generator model, achieving high precision, evidenced by an impressive AUC of $0.95$.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2024 26 February - 1 March 2024, San Diego, CA, USA ISBN 1-891562-93-2 https://dx.doi.org/10.14722/ndss.2024.23xxx www.ndss-symposium.org
I Introduction
The recent developments in image-generative models have been remarkably swift, and many applications based on these models have appeared. Diffusion models [41, 18, 35, 40, 46, 53, 56, 44, 19, 54] have come to the forefront of image generation. These models generate target images by progressive denoising a noisy sample from an isotropic Gaussian distribution. In an effort to expedite the training of diffusion models and reduce training expenses, Stable Diffusion [44] was introduced. Leveraging the extensive and high-fidelity LAION-2B [48] dataset for training, the Stable Diffusion pre-trained checkpoint, available on HuggingFace, can be fine-tuned efficiently with just a few steps for effective deployment in downstream tasks. This model’s efficiency has spurred an increasing number of usages of Stable Diffusion.
@@ I Introduction
In a white-box attack scenario, attackers have access to all parameters of a model. Similar to membership inference attack targeting classifiers, attacks against diffusion models also utilize internal model information such as loss [6, 20, 34] or gradients [37] as attack features. Hu et al. [20] and Matsumoto et al. [34] have utilized losses at different timesteps of querying the model as attack features. Similarly, Carlini et al. [6] employed losses across various timesteps but incorporated the LiRA framework to construct two distributions for inferring the membership of a sample $x$. Pang et al. [37] took a different approach by using the model’s gradients at different timesteps as the attack features, positing that gradient information better reflects the model’s response to $x$.
@@
Although white-box attacks can achieve high success rates, their limitation lies in the requirement for complete access to the target model’s information, which is often impractical in real-world scenarios. Compared with white-box attack, gray-box approaches do not require full access to the model’s parameters; instead, they only necessitate the intermediate outputs from the diffusion model during the denoising process to serve as features for inference [12, 14, 20, 25]. For example, Duan et al. [12], and Kong et al. [25] have leveraged the deterministic nature of DDIMs, using the approximated posterior estimation error of intermediate outputs at different timesteps as attack features. Hu et al. [20] have proposed using intermediate outputs to estimate the log-likelihood of samples as attack features. However, these attacks inevitably rely on the intermediate images generated during the model’s operation. In real-world scenarios, if a malicious model is trained using private or unsafe images, typically only the final output image is provided, with efforts made to conceal as many model details as possible. Therefore, the more practical scenarios would be black-box.
@@
There are also black-box attacks for GANs [7, 17] and VAEs [17]. These are based on unconditional generative models and involve a highly stochastic generation process that requires extensive sampling for inference, which becomes inefficient when directly applied to diffusion models. The other black-box attacks [34, 61, 63], although more tailored for diffusion models, focus on simulations and lack the necessary conditions to be used in realistic scenarios. We will discuss them in Section II-E.
@@
In this paper, we present a black-box attack framework fit for state-of-the-art image-generative models. The framework was built on a careful analysis of the objective function of the diffusion model as its theoretical foundation. It also incorporates four potential attack scenarios tailored for different settings of diffusion models. We demonstrate the efficacy of our attack using the pre-trained Stable Diffusion v1-5 and further validate it fine-tuned with CelebA-Dialog [23], WIT [57], and MS COCO datasets [30].
Compared with existing black-box methods [34, 63], our attack under four attack scenarios can get $87\%$ accuracy and outperform other methods by nearly $35\%$. We systematically evaluate all components, including the image encoder, distance metrics, inference steps, and training set sizes. Our method is able to achieve high AUC scores across three datasets: $0.95$, $0.85$, and $0.93$. Even using different types of generative models as shadow models to employ the attack, our attack still can get at least $83\%$ for four attack scenarios on three datasets. The results show that our attack is robust and fit for real-world requirements. To further comprehensively evaluate our attack, we employed DP-SGD [1] as a defensive strategy to assess the attack’s effectiveness. By reducing the model’s ability to memorize training samples, DP-SGD defends against our attack. This finding is consistent with the outcomes observed in other attacks [6, 20, 12].
@@
  • Consistent with the definition in Suya et al. [58], four attack scenarios are considered in which an attacker can perform an attack based on the query access as well as the quality of the initial auxiliary data, and three different attack models are used to determine the success rate of the attack, respectively.
  • The efficacy of the attack is evaluated on the CelebA, WIT, and MS COCO datasets using fine-tuned Stable Diffusion v1-5 as the representative target model. The attack’s impact is analyzed by considering various factors: image encoder selection, distance metrics, fine-tuning steps, inference step count, member set size, shadow model selection, and the elimination of fine-tuning in the captioning model.
@@
Section II reviews key works on denoising generative models and membership inference attacks, including their application against diffusion models. Section III introduces our scores-based black-box attack on diffusion models, tailored to four levels of attacker knowledge. Section IV describes our experimental setup, and Section V compares our attack’s effectiveness with existing methods and examines various influencing factors. Section VI shows that our attacks remain effective under common defenses. Section VII discusses some other research related to our work. Section VIII concludes the paper, summarizing our main findings and contributions.
II Background
II-A Machine Learning
@@
II-A1 Classification Models
In the context of classification model training, the objective is to map an input $x$ to a category $y$. The functional representation of the model can be expressed as $y=\mathcal{M}(x)$, where $x$ denotes the input (e.g., an image), $\mathcal{M}$ represents the classification model, and $y$ denotes the corresponding label. The loss in the classification model, which quantifies the discrepancy between the predicted and true labels, can be articulated as follows:
@@ II-A1 Classification Models
II-A2 Generative Models
Generative models are designed to generate $\hat{x}=\mathcal{G}(z)$, where $z$ is the randomness not provided by users but inherent to the server hosting the generator $\mathcal{G}$.
Popular generative models include VAEs [24], GANs [15], and diffusion models [18]. Recently, diffusion models have gained significant traction. Building on the classical DDPM (Denoising Diffusion Probabilistic Models), a plethora of models, such as Imagen [46], DALL·E 3 [3], GLIDE [35], Stable Diffusion [44], have emerged and can generate high-quality images based on prompt information. In this paper, we mainly focus on diffusion models.
II-B Diffusion Models
II-B1 Foundation of Diffusion Models
@@
In Equation 1, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, where $\alpha_{i}$ is a predefined parameter that decreases incrementally within the interval $[0,1]$. The term $\epsilon_{t}$ is a random Gaussian noise derived using the reparameterization trick from multiple previous forward steps (more details in Appendix A).
@@
II-B2 Prompt Guided Diffusion Models
Diffusion models [3, 40, 46, 35, 44] mentioned above are also capable of generating high-quality images conditional on prompt information $p$, denoted as $\hat{x}=\mathcal{G}(z,p)$ (further details can be found in Appendix B). Our experiments primarily utilize the current publicly available state-of-the-art model, Stable Diffusion [44]. Distinct from other diffusion generative models [35, 40, 46], Stable Diffusion uniquely conducts both the forward and reverse processes within the latent space. This approach offers advantages: the noise addition and removal processes operate over a smaller dimensionality, allowing for faster model training at lower computing costs. Additionally, within the latent space, the model can accommodate diverse prompt information to guide image generation. Importantly, Stable Diffusion is open-sourced and provides multiple high-quality pre-trained checkpoints online. This aligns well with the focus of our study on potential privacy concerns when fine-tuning pre-trained models for downstream tasks.
II-C Membership Inference Attacks
@@
To formalize membership inference attacks, assume there is a data sample $x$, a model $\mathcal{M}_{\theta}$ trained with dataset $\mathcal{D}$. The attack $\mathcal{A}$ will access $\mathcal{M}_{\theta}$, $\mathcal{D}$ and take data sample $x$ as input, the output a bit $b\leftarrow\mathcal{A}^{\mathcal{D}}(x,\mathcal{M}_{\theta})\in\{0,1\}$ indicating whether $x$ was used in training (i.e., $x\in\mathcal{D}$) or not. For simplicity, we use $\theta$ denoted model $\mathcal{M}_{\theta}$ and omit $\mathcal{D}$.
@@
Carlini et al. [5] suggest the use of confidence scores as attack features and the creation of two distributions, $\mathbb{D}_{\text{in}}$ and $\mathbb{D}_{\text{out}}$, based on the confidence scores of samples from the member and non-member sets, respectively. The distributions are then utilized to calculate the probability density function of query data $x$ in the member set and non-member set.
@@
MIAs against Diffusion Models.
@@
In the context of MIA against diffusion models, due to the structural differences between diffusion models and classification models, as well as the dissimilarities in their inputs and outputs, MIAs designed for classification models cannot be directly applied to diffusion models. The focus of research lies in how to construct features for MIA. We classify existing attacks against diffusion models as white-box, gray-box, and black-box, and introduce them separately. In white-box attacks, methods in this setting exploit the loss (derived from each timestep using Equation 2) and gradients (via backpropagation through the model). Gray-box attacks typically necessitate access to a model’s intermediate outputs but do not require any internal model information. For gray-box attacks targeting diffusion models, the model’s denoising trajectory, particularly the noisy images, is utilized as attack data. In contrast, black-box attacks operate without knowledge of the model’s internal mechanics or process outputs, relying solely on the final generated images for analysis. In Table I, we compare all existing attacks. Each type of attack’s details are deferred to Section II-E1 and Section VII.
II-D Problem Formulation
II-D1 Threat Model
@@
Given the target sample $x$ and black-box access to the target image-generative model $\mathcal{G}$, the goal of the attacker is to determine whether $x$ was used to train $\mathcal{G}$. More specifically, we focus on the fine-tuning process, namely, we care about the privacy of the fine-tuning dataset of $\mathcal{G}$, and do not care about the pre-training dataset. We focus on fine-tuning because (1) the attacks will be similar for direct training, while the computational cost for experiments on fine-tuning MIA will be much smaller, and (2) the pretraining and fine-tuning paradigm is more popular with modern large models, and we will provide motivations for focusing on the fine-tuning process later.
@@
We categorize the threat model into four scenarios (as shown in Table I) with two dimensions, namely:
@@
  • Target Sample Component. One distinct property of the current image generator models is that there exists the flexibility to input text prompt to guide model generation. Two configurations for the attacker can be considered: First, the query data $x$ aligns with the training data as a text-image pair ($x=\langle T_{q},I_{q}\rangle$, where $T_{q}$ denotes the text component and $I_{q}$ denotes the corresponding image component). Second, the attacker only obtains a suspect image potentially revealing private information without a corresponding caption ($x=I_{q}$). As our focus here is MIAs on text-to-image generative models, the scenario where $x$ solely consists of text is not deemed practical and hence, is not discussed.
@@
  • Auxiliary Dataset. Similar to all other MIAs, we assume an auxiliary dataset $\mathbf{D}^{\prime}$ is available. It is used to train the shadow models $\mathcal{G}^{s}$ to mimic the behavior of the target model. We consider two scenarios for $\mathcal{D^{\prime}}$: $\mathbf{D}^{p}$ and $\mathbf{D}^{s}$, indicating whether the attacker has access to real samples used to fine-tune $\mathcal{G}$ or just samples from the same distribution.
@@
II-D2 Motivation
Our work aims to reveal privacy leakages of the fine-tuning dataset when fine-tuning a pre-trained image generative model. We focus on the fine-tuning phase because, from the model publishers’ perspective, training a generative model from scratch would entail significant computational costs, and the most efficient strategy for them would be to fine-tune a model that has already been pre-trained on a large-scale dataset. Given the full open-source nature of the Stable Diffusion [44], and the extensive availability of pre-trained models capable of generating photorealistic images from entities like CompVis [^2] https://huggingface.co/CompVis/stable-diffusion and Stability AI [^3] https://github.com/Stability-AI/generative-models, there has been an increasing trend of leveraging these pre-trained models for fine-tuning to specific downstream tasks. Furthermore, an increasing number of companies, such as Amazon [^4] https://aws.amazon.com/sagemaker/jumpstart/, OctoML [^5] https://octoml.ai/blog/the-beginners-guide-to-fine-tuning-stable-diffusion/, and CoreWeave [^6] https://docs.coreweave.com/cloud-tools/argo, are offering services in this domain. The data privacy issue during the training of these downstream tasks has not been explicitly studied. Our work seeks to uncover privacy leakage issues in this process and raise awareness of them.
@@ II-D2 Motivation
II-E Existing Solutions
II-E1 Black-box MIA against Traditional Image-generatve Models
@@
GAN-Leaks [7] employs a similar intuition and, instead of density in the Monte-Carlo attack [17]), roughly speaking, use the shortest distance of the generated samples from the target sample as the criteria. It also proposes another attack assuming an extra ability to optimize the noise input $z$ to the generator (which is not strictly the black-box setting; we will describe it and compare with it in the evaluation) so it can reduce the number of generated samples. More formal details about these attacks are deferred to Appendix D.
The reason we cannot apply Monte-Carlo attack [17] and GAN-Leaks [7] to diffusion models is that both attack methods require the model to sample a large number of images. Diffusion models progressively denoise during the inference process, involving dozens of steps, unlike VAEs and GANs, which require only a single step. Both Monte-Carlo attack and GAN-Leaks need to construct $100K$ samples to achieve optimal attack performance [7]. For the diffusion model, this will take even hundreds of times longer in terms of computing time. Furthermore, these attacks are unsuitable for conditional generative models. Although GAN-Leaks proposed a partial-black attack, conditional embedding (e.g., text embedding) in diffusion models is significantly more complex than the initial noise $z$ in GANs and VAEs. Therefore, traditional black-box MIAs are not feasible for current diffusion models.
II-E2 Black-box MIA against Recent Diffusion Models